OpenStudy (anonymous):

FAN AND MEDAL!!! Solve for x: 2|x - 3| + 1 = 7 I got x = 5.5 and x = 1.5

OpenStudy (anonymous):

x = 1.5 is wrong

OpenStudy (anonymous):

there both wrong

OpenStudy (anonymous):

x=0

OpenStudy (anonymous):

can you show your steps?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

mine?

OpenStudy (anonymous):

2|x - 3| + 1 = 7 2|x - 3|=6

OpenStudy (anonymous):

|x - 3|=3

OpenStudy (anonymous):

x=6 or 0

OpenStudy (anonymous):

hold on! I see where I made a mistake!

OpenStudy (anonymous):

There should only be one x answer

OpenStudy (anonymous):

lol there should be 2

OpenStudy (anonymous):

6 is work

OpenStudy (anonymous):

its absolute value, theres always 2 or no answer

OpenStudy (anonymous):

Hmm, I was thinking that was square root

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

What did you end up for your answers?

OpenStudy (anonymous):

I got x = 9/2 and x=-9/2

OpenStudy (anonymous):

now I got x = 5 and x = 0....but that's wrong and so it @jtryon

OpenStudy (anonymous):

*is

OpenStudy (anonymous):

Oh, I see what I did

OpenStudy (anonymous):

A) x = 0, x = 6 B) x = -1, x = 6 C) x = 0, x = 9 D) no solutions

OpenStudy (anonymous):